OpenStudy (anonymous):

Simplify square root of 2 multiplied by the cube root of 2. A: 2 to the power of 1 over 6 B: 2 to the power of 2 over 3 C: 2 to the power of 5 over 6 D: 2 to the power of 7 over 6

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@amistre64

OpenStudy (tkhunny):

Hint: \(\sqrt{2}\cdot\sqrt[3]{2} = 2^{1/2}\cdot 2^{1/3}\)

OpenStudy (amistre64):

you will simply need to memorize properties of exponents then, they should be in your notes or material

OpenStudy (amistre64):

what happens to exponents when we multiply like bases?

OpenStudy (anonymous):

You add them?

OpenStudy (amistre64):

we can demonstrate the process with integer exponents b^2 * b^3 is a good example, lets expand them (bb)(bbb) now the paranthesis are really pointless so what we have is bbbbb, and converting back to an exponent is just b^5 2 bs plus 3bs is 5 bs

OpenStudy (amistre64):

yes, you add the exponents when mutliplying like bases :)

OpenStudy (anonymous):

Honestly, I have no idea how to do this and it makes me feel so stupid.

OpenStudy (anonymous):

And I'm so stressed with the fact of having to get caught up that I can't even concentrate.

OpenStudy (amistre64):

playing catchup is never fun .... but thats the hand we are dealt at times. good luck

OpenStudy (anonymous):

I just need help on this one question! I already did all the others, this was a 50 question quiz..